# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v2

@@ 5 Implementation
We employ the HashGrid [12] for $\tau$, although any encoding using trilinear interpolation we defined can be applied, while our discussion can be found in Section B.2. We illustrate our method in Algorithm 1 and 2, describing the details in the following sections.
@@
While the original edge subdivision algorithm [4] starts with a unit cube of eight vertices and a dozen edges, we know that the unit cube is subdivided by orthogonal planes consisting of the grid. Let an input coordinate be ${\bm{x}}\in[0,1]^{3}$, and there is a unit cube such that one of its corners is at the origin. Taking into account multi-resolution grids, we obtain the marks $m_{i}$, such that the grid planes are $x=m_{i}$, $y=m_{i}$, and $z=m_{i}$, following Algorithm 4 in Appendix. Notice that we carefully consider the offset $s/2$, which prevents the zero derivatives from aligning across all resolution levels (ref. Appendix A of Müller et al. [12]). Leveraging the obtained marks, we derive the associated grid vertices and edges, which serve as the initial sets for ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively.
@@
The number of orthogonal planes may significantly surpass the number of neurons, making it impractical to allocate elements for the sign-vectors (Definition 3.4). To address this, leveraging the orthogonality of the grid planes, we store a plane index along with an indicator of whether the input is on the plane ($\gamma=0$) or not ($\gamma=1$), similarly.
@@ 6 Experiment
For a given trilinear neural network with the eikonal constraint, our goal is to extract meshes from it in a way that accurately reflects the learned SDF. (Note that NeuS [6] showed how to convert the density networks of NeRFs to an SDF in the frameworks of volume rendering.) To evaluate this, we utilize marching cubes with excessive sampling to get a pseudo-ground truth mesh.
@@
For ReLU neural networks, the number of layers $L$ of 3, hidden size $H$ of 16, and $\epsilon$ of 1e-4 for the sign-vectors (ref. Definition 3.4). The weight for the eikonal loss is 1e-2. For HashGrid, the resolution levels of 4, feature size of 2, base resolution $N_{\text{min}}$ of 2, and max resolution $N_{\text{max}}$ of 32 by default. $N_{\text{min}}$ and $N_{\text{max}}$ are doubled (x2) or quadrupled (x4) for Medium or Large settings in Figure 3. We use the official Python package of tinycudnn  [^4] https://github.com/NVlabs/tiny-cuda-nn for the HashGrid module.
Chamfer distance.
@@
$\text{CD}({\mathcal{S}}_{0},{\mathcal{S}}_{1})$ $=\frac{1}{2}\Big{(}\overrightarrow{\text{CD}}({\mathcal{S}}_{0},{\mathcal{S}}_{1})+\overrightarrow{\text{CD}}({\mathcal{S}}_{1},{\mathcal{S}}_{0})\Big{)},$ (10)
$\overrightarrow{\text{CD}}({\mathcal{S}}_{0},{\mathcal{S}}_{1})$ $=\frac{1}{|{\mathcal{S}}_{0}|}\sum_{{\bm{x}}\in{\mathcal{S}}_{0}}\min_{{\bm{y}}\in{\mathcal{S}}_{1}}\|{\bm{x}}-{\bm{y}}\|_{2}^{2}.$ (11)
@@
Table 1 and Figure 3 show the results for the Standford 3D Scanning repository [30]. Varying the number of samples in marching cubes [9], we plot the baseline with respect to the number of generated vertices. Given that our method extracts vertices from the intersection points of hypersurfaces, it enables parsimonious representations using the lower number of vertices having better CD, summarized by chamfer efficiency (CE) in Table 1, and ours are located in the lower-left region compared with the baselines in Figure 3. In particular, we observe a strong CE in a simple object, e.g., Drill Bit, which achieved a CE of 47.8 versus 19.2 (MC).
Our method inherently holds an advantage over MC in capturing optimal vertices and faces, especially in cases where the target surfaces exhibit planar characteristics. When dealing with curved surfaces, the smaller grid size of the Large model in Figure 3 views a curved surface in a smaller interval, and it tends to approach the plane. From the plotting of Small to Large, our CDs (colored crosses) are decreasing to zero, consistently with better CEs, confirming this speculation. The complete results of the Large models for the Stanford 3D Scanning repository can be found in Tables 6, 7 and 8 in Appendix F and its visualization can be found in the right side of Figure 5.
Table 1: Chamfer distance (CD) and chamfer efficiency (CE) for the Stanford 3D Scanning repository [30]. The CD ($\times$1e-6) is evaluated using the marching cubes (MC) with 2563 grid samples as the reference ground truth for all measurements. The CE is defined as $100/(\sqrt{|{\mathcal{V}}|\times\text{CD}})$, providing a concise representation of the trade-off between the number of vertices and the CD. Please refer to Table 2 and Table 3 for the full results, and Table 4 for standard deviation and time spent in Appendix (Ours took 0.97 $\pm$ 0.21 while MC also took within 1 sec for the bunny.)
@@
To assess the efficacy of the eikonal constraint, which enforces the planarity of hypersurfaces as outlined in Theorem 4.5, we quantify the flatness error associated with the plenary constraints presented in the proof of Theorem D.5. Let ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ be two vertices consisting of a diagonal edge, while ${\bm{x}}_{1\dots 6}$ are its remaining cubic corners. We obtain $\tilde{\nu}({\bm{x}}_{1\dots 6})$ in a piecewise trilinear region as described in Section 5. This evaluation is conducted as follows:
$\Delta_{\text{flat}}=$ $\mathbb{E}_{\mathcal{E}}\Big{[}\frac{1}{2}\big{(}\|\Sigma_{i\in\{1,2,4\}}\tilde{\nu}({\bm{x}}_{i})\|_{1}+\|\Sigma_{i\in\{3,5,6\}}\tilde{\nu}({\bm{x}}_{i})\|_{1}\big{)}$
$~{}~{}+\frac{1}{3}\Big{(}\sum_{i=1}^{3}\|\Sigma_{j\in\{i,7-i\}}\tilde{\nu}({\bm{x}}_{j})\|_{1}\Big{)}\Big{]}/2.$ (13)
@@
We provide the qualitative analyses in Figures 5, 6, 7 and 8 in Appendix G encompassing the comparison between the Small and Large models (Figure 5), visualizing the marching cubes varying the number of samples (Figure 6), detailed comparisons (Figure 7), and the gallery of the Stanford 3D Scanning meshes from the Large models (Figure 8).
@@
Although we implement the algorithm using pre-defined parallel tensor operations in PyTorch [31], achieving time and memory complexity of $\mathcal{O}(|\mathcal{V}|)$ as detailed in Appendix E, our method inherently requires forward passing the intermediate vertices during edge subdivisions iterating over neurons, which depends on previous steps (Algorithm 1). For the Stanford bunny using the Small model, the intermediate counts for vertices and edges exceed 11K and 19K, respectively, although they are eventually reduced to 4.3K vertices for face extraction. This overhead is negligible for the Small models, taking 0.97 seconds; however, for the Large models, the process takes 69.0 seconds for 136K vertices, compared to 11.88 seconds for marching cubes with 512 samples per axis. We note that parallelizing over multiple GPUs would be a reliable option to improve latency.
@@ 7 Conclusions
In conclusion, our exploration of novel techniques in visualizing trilinear neural networks unveils intricate structures and extends mesh extraction applications beyond CPWA functions. Focused on trilinear interpolating methods, our study yields novel theoretical insights and a practical methodology for precise mesh extraction. Notably, Theorem 4.5 shows the transformation of hypersurfaces to planes within the trilinear region under the eikonal constraint. Empirical validation using chamfer distance and efficiency, and angular distance affirms the correctness and parsimony of our methodology. The correlation analysis between eikonal loss and hypersurface planarity confirms the theoretical findings. Our proposed method for approximating intersecting points among three hypersurfaces, coupled with a vertex arrangement methodology, broadens the applications of our techniques. This work establishes a foundation for future research and practical applications in the evolving landscape of deep neural network analysis and toward real-time rendering mesh extraction.
@@
I would like to express my sincere appreciation to my brilliant colleagues, Sangdoo Yun, Dongyoon Han, and Injae Kim, for their contributions to this work. Their constructive feedback and guidance have been instrumental in shaping the work. The NAVER Smart Machine Learning (NSML) platform [32] had been used for experiments.
@@ A.1 Tropical geometry
The tropical semiring is an algebraic structure that extends real numbers and $\infty$- ∞ with the two operations of tropical sum $\oplus$ and tropical product $\odot$. Tropical sum and product are conveniently replaced with $\max(\cdot)$ and $+$, respectively. Notice that $\infty$- ∞ is the tropical sum identity, zero is the tropical product identity, and these satisfy associativity, commutativity, and distributivity. For example, a classical polynomial $x^{3}+2xy+y^{4}$ would represent $\max(3x,~{}x+y+2,~{}4y)$. Roughly speaking, this is the tropicalization of the polynomial, denoted tropical polynomial or simply $f$.
@@ A.2 Tropical hypersurface
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$. In other words, this is the set of points where $f({\bm{x}})$ equals two or more monomials in $f$.
@@ A.3 Tropical algebra of neural networks
Without loss of generality, $L$-layer neural networks with ReLU activation are a composition of functions defined as:
@@
Proposition A.3.
(Decision boundary of neural networks, restated) The decision boundary for $L$-layer neural networks with ReLU activation $\nu^{(L)}({\bm{x}})=0$ is where two tropical monomials, or two arguments of $\max$, equal (Definition 3.1), recursively, as follows:
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=H_{j}^{(i)}({\bm{x}})-G_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)\big{\}}$
where the subscription $j$ denotes the $j$-th element, assuming the hidden size of neural networks is $H$, while the $L$-th layer has a single output for rather discussions as an SDF. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
This implies that piecewise hyperplanes corresponding to the preactivations $\{\nu_{j}^{(i)}\}$ are the candidates shaping the decision boundary. Each linear region has the same pattern of zero masking by $\textrm{ReLU}(\cdot)$, effectively making it linear as an equation of a plane. In turn, if the pattern changes as an input is continually changed, it moves to another linear region. From a geometric point of view, this is why the number of linear regions grows polynomially with the hidden size $H$ and exponentially with the number of layers $L$ [27, 18].
Appendix B Further discussions
B.1 Satisfying the eikonal equation and approximation
Theorem 4.7 handles where the eikonal equation does not make perfect hyperplanes. First, we observed the characteristics of trilinear hypersurfaces as illustrated in Figure 4 in Appendix. Among 64 edge scenarios within a trilinear region, a diagonal plane intersects with the hypersurface in nearly all cases. Moreover, we can take advantage of it since we can eliminate one variable by assuming the curved edge lies on the $x=z$ plane. In this light, Theorem 4.7 used linear algebra to rearrange two trilinear equations to get Equations 77 and 74. As we mentioned in the manuscript, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error. This error is tolerated by the hyperparameter $\epsilon$=1e-4 (Section 6) as specified in Definition 3.4.
B.2 On the claim that any encoding using trilinear interpolation can be applied
In this work, we describe “any encoding using trilinear interpolation (that) we defined can be applied.” The definition is implicitly referred to Definition 4.1 (Trilinear interpolation), where the cubic corners are represented by ${\bm{H}}\in\mathbb{R}^{D\times F}$. Since we do not rely on a specific mapping function $f:\mathrm{x}\rightarrow({\bm{H}}_{0},{\bm{H}}_{1}\dots\bm{H}_{7})$, we may apply our extraction method to any encoding using trilinear interpolation that we defined. Specifically, InstantNGP [12] used a hashing function to map a corner point to the arbitrary entry of hash tables, allowing gracious collisions with online adaptivity (refer to their Sec. 3), while TensoRF [13] can be interpreted that it used a mapping function to the summation of the multiplications of two vectors from the vector-matrix factorization. Therefore, if we have the full ranks for $R_{1},R_{2},R_{3}$, in Eqn. 3 in Chen et al. [13], among the VM factorization, which can be equivalently represented by the HashGrid with a sufficient number of hash table sizes.
B.3 Zero-trailing binarization function

### v3

@@ 5 Implementation
We employ the HashGrid [12] for $\tau$, while our discussion on its generalizability is provided in Section C.2. We illustrate our method in Algorithm 1, 2, and Figure 2, inspired by Hanin and Rolnick [30] for its visualization of linear regions. We describe the implemental details of the algorithms in the following sections and the corresponding caption of Figure 6.
@@
While the original edge subdivision algorithm [4] starts with a unit cube of eight vertices and a dozen edges, we know that the unit cube is subdivided by orthogonal planes consisting of the grid. Let an input coordinate be ${\bm{x}}\in[0,1]^{3}$, and there is a unit cube such that one of its corners is at the origin. Taking into account multi-resolution grids, we obtain the marks $m_{i}$, such that the grid planes are $x=m_{i}$, $y=m_{i}$, and $z=m_{i}$, following Algorithm 4 in Appendix. Notice that we carefully consider the offset $s/2$, which prevents the zero derivatives from aligning across all resolution levels (ref. Appendix A of Müller et al. [12]). Leveraging the obtained marks, we derive the associated grid vertices and edges, which serve as the initial sets for ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively. Please refer to Figure 2(a) where its multi-resolution grids are visualized using randomized colors.
@@
The number of orthogonal planes may significantly surpass the number of neurons, making it impractical to allocate elements for the sign-vectors (Definition 3.4). To address this, leveraging the orthogonality of the grid planes, we store a plane index along with an indicator of whether the input is on the plane ($\gamma=0$) or not ($\gamma=1$).
@@ 6 Experiment
For a given trilinear neural network with the eikonal constraint, our goal is to get a boundary mesh of the zero-set of the SDF. (Note that NeuS [6] showed how to convert the density networks of NeRFs to an SDF in the frameworks of volume rendering.) To evaluate this, we utilize marching cubes with excessive sampling to get a pseudo-ground truth mesh to remove Bayes error caused by underfitting. We explicitly specify the samplings in the results.
@@
We used the number of layers $L$ of 3 and hidden size $H$ of 16 for the networks, and $\epsilon$ of 1e-4 for the sign-vectors (ref. Definition 3.4). The weight for the eikonal loss is 1e-2. For HashGrid, the resolution levels of 4, feature size of 2, base resolution $N_{\text{min}}$ of 2, and max resolution $N_{\text{max}}$ of 32 by default. $N_{\text{min}}$ and $N_{\text{max}}$ are doubled (x2) or quadrupled (x4) for Medium or Large settings in Figure 5. We use the official Python package of tinycudnn  [^4] https://github.com/NVlabs/tiny-cuda-nn for the HashGrid module.
Chamfer distance and efficiency.
@@
$\text{CD}({\mathcal{S}}_{0},{\mathcal{S}}_{1})$ $=\frac{1}{2}\Big{(}\overrightarrow{\text{CD}}({\mathcal{S}}_{0},{\mathcal{S}}_{1})+\overrightarrow{\text{CD}}({\mathcal{S}}_{1},{\mathcal{S}}_{0})\Big{)},~{}~{}\overrightarrow{\text{CD}}({\mathcal{S}}_{0},{\mathcal{S}}_{1})=\frac{1}{|{\mathcal{S}}_{0}|}\sum_{{\bm{x}}\in{\mathcal{S}}_{0}}\min_{{\bm{y}}\in{\mathcal{S}}_{1}}\|{\bm{x}}-{\bm{y}}\|_{2}^{2}.$ (9)
@@
Figure 3: Chamfer distance for the bunny with the Large model comparing with MC, MT, and NDC.
Table 1 and Figure 5 show the results for the Standford 3D Scanning repository [31]. Varying the number of samples in marching cubes [9], we plot the baseline with respect to the number of generated vertices. Given that our method extracts vertices from the intersection points of hypersurfaces, it enables parsimonious representations using the lower number of vertices having better CD, summarized by chamfer efficiency (CE)  [^5] CE is defined as $100/(\sqrt{|{\mathcal{V}}|\times\text{CD}})$, reflecting the trade-off between the number of vertices and the CD. in Table 1, and ours are located in the lower-left region compared with the baselines in Figure 5. In particular, we observe a strong CE in a simple object, e.g., Drill Bit, which achieved a CE of 47.8 versus 19.2 (MC).
Our method inherently holds an advantage over MC in capturing optimal vertices and faces, especially in cases where the target surfaces exhibit planar characteristics. When dealing with curved surfaces, the smaller grid size of the Large model in Figure 5 views a curved surface in a smaller interval, and it tends to approach the plane. From the plotting of Small to Large, our CDs (colored crosses) are decreasing to zero, consistently with better CEs, confirming this speculation. The complete results of the Large models and its visualization can be found in Tables 6, 7 and 8 in Section F.1 and the right side of Figure 10 in Appendix, respectively. We also compare ours with Marching Tetrahedra (MT) [16, 10] and Neural Dual Contour (NDC) [32] in Figure 3, where our method achieved the lowest chamfer distance and the most efficiency method with respect to the number of vertices. For the rationale behind our selection, please refer to Section F.3.
Table 1: Chamfer distance (CD) and chamfer efficiency (CE) for the Stanford 3D Scanning repository [31]. The CD ($\times$1e-6) is evaluated using the marching cubes (MC) with 2563 grid samples as the reference ground truth for all measurements. Please refer to Table 2 and Table 3 for the full results, and Table 4 for standard deviation and time spent in Appendix (Ours took 0.97 $\pm$ 0.21 while MC also took within 1 sec for the bunny.)
@@
To assess the efficacy of the eikonal constraint, which enforces the planarity of hypersurfaces as outlined in Theorem 4.5, we quantify the flatness error associated with the planarity constraints presented in the proof of Theorem D.5. Let ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ be two vertices consisting of a diagonal edge, while ${\bm{x}}_{1\dots 6}$ are its remaining cubic corners. We obtain $\tilde{\nu}({\bm{x}}_{1\dots 6})$ in a piecewise trilinear region as described in Section 5. This evaluation is conducted as follows:
$\Delta_{\text{flat}}=$ $\mathbb{E}_{\mathcal{E}}\Big{[}\frac{1}{4}\big{(}\|\Sigma_{i\in\{1,2,4\}}\tilde{\nu}({\bm{x}}_{i})\|_{1}+\|\Sigma_{i\in\{3,5,6\}}\tilde{\nu}({\bm{x}}_{i})\|_{1}\big{)}+\frac{1}{6}\big{(}\sum_{i=1}^{3}\|\Sigma_{j\in\{i,7-i\}}\tilde{\nu}({\bm{x}}_{j})\|_{1}\big{)}\Big{]}.$ (11)
@@
We provide the qualitative analyses in Figure 6 comparing with MT [16, 10] and NDC [32], along with MC. While the competitors exhibit over-smoothing or inaccuracies, particularly in the nose region, our method preserves surface details with consistent normals and outperforms in both accuracy and generalization. In addition, we present Figures 10, 11, 12 and 13 in Appendix G encompassing the comparison between the Small and Large models (Figure 10), visualizing the marching cubes varying the number of samples (Figure 11), detailed comparisons (Figure 12), and the gallery of the Stanford 3D Scanning meshes from the Large models (Figure 13).
Figure 6: This figure provides a detailed visualization of Figure 10 in the Appendix and its competitors. MC 64, MT 32 (Marching Tetrahedra), and NDC (Neural Dual Contour [32]) suffer over-smoothing or inaccuracy of the actual surface in the Small networks within the nose, whereas our method reflects these details (see the boundaries of a nose) with consistent normals (in colors). MT efficiently demands SDF values by utilizing six tetrahedra within grids to get intermediate vertices. However, it is inefficient in terms of the number of vertices extracted. NDC faces a generalization issue for a zero-shot setting for the given networks, producing unsmooth surfaces.
@@
Although we implement the algorithm using pre-defined parallel tensor operations in PyTorch [33], achieving time and memory complexity of $\mathcal{O}(|\mathcal{V}|)$ as detailed in Appendix E, our method inherently requires forward passing the intermediate vertices during edge subdivisions iterating over neurons, which depends on previous steps (Algorithm 1). For the Stanford bunny using the Small model, the intermediate counts for vertices and edges exceed 11K and 19K, respectively, although they are eventually reduced to 4.3K vertices for face extraction. This overhead is negligible for the Small models, taking 0.97 seconds; however, for the Large models, the process takes 4.15$\pm$0.61 seconds for 140.3$\pm$5.6K vertices with our optimized algorithm, compared to 1.31 and 11.88 seconds for marching cubes with 256 and 512 samples per axis, respectively. We note that parallelizing over multiple GPUs would be a reliable option to improve latency. For the discussions on satisfying the eikonal equation in practice and the generalization to other positional encodings using trilinear interpolation, please refer to Section C.1 and Section C.2, respectively.
@@ 7 Conclusions
In conclusion, our exploration of novel techniques in visualizing trilinear neural networks unveils its intricate structures and extends mesh extraction applications beyond CPWA functions. Focused on trilinear interpolating methods, our study yields novel theoretical insights and a practical methodology for precise mesh extraction. Notably, Theorem 4.5 shows the transformation of hypersurfaces to planes within the trilinear region under the eikonal constraint. Empirical validation using chamfer distance and efficiency, and angular distance affirms the correctness and parsimony of our methodology. The correlation analysis between eikonal loss and hypersurface planarity confirms the theoretical findings. Our proposed method for approximating intersecting points among three hypersurfaces, coupled with a vertex arrangement methodology, broadens the applications of our techniques. This work establishes a foundation for future research and practical applications in the evolving landscape of deep neural network analysis and toward real-time rendering mesh extraction.
@@
I would like to express my sincere appreciation to my brilliant colleagues, Sangdoo Yun, Dongyoon Han, and Injae Kim, for their contributions to this work. Their constructive feedback and guidance have been instrumental in shaping the work. I also express my sincere thanks to the anonymous reviewers for their help in improving the manuscript. The NAVER Smart Machine Learning (NSML) platform [34] had been used for experiments.
Appendix
[…]
@@ A.1 Tropical geometry
The tropical semiring is an algebraic structure that extends real numbers and $\infty$- ∞ with the two operations of tropical sum $\oplus$ and tropical product $\odot$. Tropical sum and product are conveniently replaced with $\max(\cdot)$ and $+$, respectively. Notice that $\infty$- ∞ is the tropical sum identity, zero is the tropical product identity, and these satisfy associativity, commutativity, and distributivity. For example, a classical polynomial $x^{3}+2xy+y^{4}$ would represent $\max(3x,~{}x+y+2,~{}4y)$. The classical polynomial is rewritten for a tropical polynomial by naively replacing + with $\oplus$ and $\cdot$ with $\odot$, respectively. After this, we rewrite by the definitions of tropical operations as $\max(3x,~{}x+y+2,~{}4y)$. Roughly speaking, this is the tropicalization of the polynomial, denoted tropical polynomial or simply $f$. Note that this example (ours with max notation) comes from the Wikipedia  [^6] https://en.wikipedia.org/wiki/Tropical_geometry.
Additionally, we present a graphical diagram in Figure 7 illustrating how the tropical polynomial divides the 2D space for clearer understanding. Please see the caption for further details.
Figure 7: A classical polynomial $x^{2}+y^{2}+2xy+2x+2y+2$ turns into a tropical polynomial as $\max(1+2x,1+2y,2+x+y,2+x,2+y,2)$, taking into account the implicit coefficient of 1. We redraw Figure 1 of the tropical curve from Zhang et al. [27] as a region map. Each color indicates a region where each tropical monomial has maximum.
@@ A.2 Tropical hypersurface
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c_{i}\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$ and tropical product :=$\odot$:=+⊙ := +. In other words, this is the set of points where $f({\bm{x}})$ equals two or more monomials in $f$.
@@ A.3 Tropical algebra of neural networks
Assuming input and output dimensions are consistent, $L$-layer neural networks with ReLU activation are a composition of functions defined as:
@@
Proposition A.3 (Decision boundary of neural networks, restated).
The decision boundary for $L$-layer neural networks with ReLU activation $\nu^{(L)}({\bm{x}})=0$ is where two tropical monomials, or two arguments of $\max$, equal (Definition 3.1), as follows:
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=H_{j}^{(i)}({\bm{x}})-G_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)~{}\textnormal{if}~{}i\neq L~{}\textnormal{else}~{}(1)\big{\}}$
where the subscript $j$ denotes the $j$-th neuron, assuming the hidden size of neural networks is $H$, while the $L$-th layer has a single output for rather discussions as an SDF. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
This implies that piecewise hyperplanes corresponding to the preactivations $\{\nu_{j}^{(i)}\}$ are the candidates shaping the decision boundary. Each linear region has the same pattern of zero masking by $\textrm{ReLU}(\cdot)$, effectively making it linear as an equation of a plane ${\bm{A}}{\bm{x}}+{\bm{b}}=0$ for some ${\bm{A}}$ and ${\bm{b}}$. In turn, if the masking pattern changes as an input is continually changed, it moves to another linear region. From a geometric point of view, this is why the number of linear regions grows polynomially with the hidden size $H$ and exponentially with the number of layers $L$ [27, 18].
Appendix B Implementation details
B.1 Zero-trailing binarization function